A golf ball is dropped from rest from a height of 8.10 m. It hits the pavement, then bounces back up, rising just 6.10 m before falling back down again. A boy then catches the ball when it is 1.50 m above the pavement. Ignoring air resistance, calculate the total amount of time that the ball is in the air, from drop to catch.
Answer #1

First find the time taken for the golf ball to hit ground when dropped from 8.1 m

Use formula sut +1/2at2

since the ball is dropped ,initial velocity u is zero

s=1/2gt^{2}

8.1mı = 0.5 * 9.81m/S * ti

8.1 4.905 *

t1 1.285s

==============

Now time taken to rise to 6.1 m

First we need initial velocity u for rising part

Since air resistance is ignored,kinetic energy at bottom will be equal to potential energy at height 6.1m

K Ebottom- PE6.1m

1/2mu2 = mgh

V2gh 1

u=v/2 * 9.81 * 6.1

u-10.94m/s

-----------------------

Use formula sut +1/2at2

Use formula 6.1m=10.94m/s*t_{2}-1/2gt_{2}^{2}

6.1 = 10.94 * t2-4.9056

4.905t_{2}^{2}-10.94*t_{2}+6.1=0

Solve the quadratic equation using a calculator

{color{Red} t_{2}=1.11s}

=================

Final Part-A boy then catches the ball when it is 1.50 m above the pavement

so height covered =6.1 -1.5 =4.6m

Use formula sut +1/2at2

since the ball is dropped ,initial velocity u is zero

s=1/2gt^{2}

.60.5 9.81m/s2

4.6=4.905*t_{3}^{2}

{color{Red} t_{3}=0.968s}

=========

Total time ,t=t_{1}+t_{2}+t_{3}

t=1.285s+1.11s+0.968s

ANSWER: {color{Red} t=3.363s}

====================